OpenStudy (anonymous):

Three multiplied by the sum of 4 and a number is the same as 18 more than the number. Find the number. If n is "the number," which equation could be used to solve for the number? 3n + 4 = n + 18 4(3 + n) =n + 18 3(4 + n) =n + 18 3(4 + n) + 18 > n

OpenStudy (anonymous):

Can someone explain how to do this @Hero

hero (hero):

The sum of 4 and a number is 4 + n Three times that is 3(4 + n) 18 more than a number is n + 18 3(4 + n) = n + 18

OpenStudy (anonymous):

But...wouldnt P.E.M.D.A.S apply here? so wouldnt it in fact be...wait...so... The part that states: Three multiplied by the sum of 4 and a number means that 4+n is acctualy (4+n)? so it is realy 3(4+n)=18+n

hero (hero):

PEMDAS only applies to simplifying numerical expressions. It doesn't apply to solving algebraic equations. When solving algebraic equations, you have to focus on isolating a variable..NOT PEMDAS.

OpenStudy (anonymous):

>_> so...The lesson is called two or more unknowns...so...I have to follow the order its listed in the sentence?

OpenStudy (anonymous):

Ok...That kind of makes sense hero....Can you watch me work out one more and assist me if im wrong?

OpenStudy (anonymous):

Three-sevenths of a number is 21. Find the number. is the problem...

OpenStudy (anonymous):

So...If its 3 parts of the number already...

hero (hero):

If you were solving for n, you have to isolate n.

OpenStudy (anonymous):

then i need to 21 and muliply it by 4 since it is already 3 parts of 7? and to get it to the full value id need to just multiply it by 4?

hero (hero):

\[\frac{3}{7}n = 21\]

OpenStudy (anonymous):

Ok...so instead of what ever i just attempted to say i need to take 21. divide it by 3. mulitply what i get from that by 7 and ill get n?

hero (hero):

of means "times" or "multiply"

hero (hero):

three-sevenths of a number means the same thing as three-sevenths TIMES the number.

OpenStudy (anonymous):

? so i couldnt divide 21 by 3 and mutiply the awnser from that by 7? wouldnt it give me the same thing?

OpenStudy (anonymous):

\[21 \div 3 = 7 \times 7 = 49\]

hero (hero):

Yes, you can do that. If the number on the right is divisible by the numerator, then that would be a smart way to reduce the number on the right.

OpenStudy (anonymous):

Hrmm...Could you tutor me? xD

hero (hero):

Tutor you? Just post your questions on OS. You're already being tutored...sort of.

OpenStudy (anonymous):

Ok...Well...These right here kind of...to be honest...they f--k with my head...like seriously...they make me rage... Separate 846 into 3 parts so that the second part is twice the first part and the third part is triple the second part. Which of the following equations could be used to solve the problem? 846 = x + 2x + 3x 846 = x + 2x + 6x 846 ÷ 3 = x + 2x + 6x 846 ÷ 3 = x + 2x + 3x

hero (hero):

Just do each part one at a time. Start with the smallest part and then build to the largest part.

OpenStudy (anonymous):

so...the 2nd one?

hero (hero):

The first part is the smallest part, so label it x The second part is twice the first part so 2x The third part is triple the second part so 3(2x) or 6x All three parts combined is equal to 846: x + 2x + 6x = 846

OpenStudy (anonymous):

So it is the 2nd one :D

OpenStudy (anonymous):

Ok last one hero...i want to be SURE i got this down pat. Five times a number is the same as 30 more than 8 times the number. Find the number.

OpenStudy (anonymous):

so 5 = 30+8Xn

OpenStudy (anonymous):

no wait i meant \[5 \times N = 30 + 8 \times n\]

hero (hero):

Try to avoid using the times symbol when setting up algebraic equations. Simply write \(5n = 30 + 8n\) Make sure your variables match. For example, \(N\) is not the same as \(n\)

OpenStudy (anonymous):

Ok. Understood. So...How would i solve for n? this is where i get confused...

hero (hero):

Ever heard of like terms? Put like terms on the same side.

hero (hero):

Isn't 8n and 5n like terms?

OpenStudy (anonymous):

Ok...So let me try please...

hero (hero):

The put the other terms (that are not like) on the other side.

OpenStudy (anonymous):

8N can be subtracted from both sides so the new problem would be

OpenStudy (anonymous):

\[-5N =30\]

hero (hero):

Well...let me say something about that.

OpenStudy (anonymous):

Ok

hero (hero):

What you want to do is make sure that the number that is multiplied by the variable is positive.

OpenStudy (anonymous):

Well...Wouldnt the next step be 30/-5N? so wouldnt that make it 6N?

OpenStudy (anonymous):

so N=6?

hero (hero):

\(5n = 30 + 8n\) So subtract \(5n\) from both sides: \(0 = 30 + 8n - 5n\)

hero (hero):

Now...factor out the n: \(0 = 30 + (8 - 5)n\)

hero (hero):

Now obviously 8 - 5 = 3 so: \(0 = 30 + 3n\)

hero (hero):

Subtract 30 from both sides: \(-30 = 3n\)

OpenStudy (anonymous):

so..30/3N= 10=N

hero (hero):

Now we can divide by a positive number: \(-\dfrac{30}{3} = n\) Remember to keep track of the negative and never divide by the variable.

hero (hero):

\(-10 = n\)

OpenStudy (anonymous):

-10?

hero (hero):

Are you surprised?

OpenStudy (anonymous):

Yea....a little bit...i thought that...It just...what?

hero (hero):

They didn't say the number must be positive. Negative numbers are still numbers.

OpenStudy (anonymous):

True...but...???? im trying to figure out how the 10 became negative...

OpenStudy (anonymous):

i see where it became negative but why would it have to be subtracted...couldnt you flip it to the other side flipping its sign?

OpenStudy (anonymous):

Sorry...i understand how you

OpenStudy (anonymous):

did it...thats just the way my teachers taught me...

hero (hero):

Because you want to divide only by positive numbers. Dividing by positive numbers make more sense than dividing by negative numbers.

OpenStudy (anonymous):

>_> but...OH so...ok i still use the "house" sign for dividing because im a two year old inside...so...the positive number would be on the porch and the negative is under the roof? so...your not dividing by a negative number...the negative is being divided!!!